OpenStudy (anonymous):

What is the first term of the geometric sequence whose fifth term is 1/24 and tenth term is 1/768

OpenStudy (kropot72):

\[The\ n ^{th}\ term\ is\ ar ^{n-1}\] So we can set up two equations as follows: \[\frac{1}{24}=ar ^{4}............(1)\] \[\frac{1}{768}=ar ^{9}...........(2)\] Now we need to solve these equations to find the first term a and the common ratio r. Do you follow so far?

OpenStudy (anonymous):

Let the first term be a, and the common ratio be r. Then \[1/24=a r^{4}\] and \[1/768=ar ^{9}\]. Hence you may divide to obtain\[r^{5}=ar ^{9}/ar ^{4}=\frac{ \frac{ 1 }{ 768 } }{ \frac{ 1 }{ 24 } }= \frac{ 1}{ 32 }\] which implies that r=1/2. Go back to the equation for 1/24 and solve to obtain a=2/3.

OpenStudy (tesa12345):

so whats the answer

OpenStudy (tesa12345):

@kropot72

OpenStudy (kropot72):

@tesa12345 Can you follow the solution by @francisgcm?

OpenStudy (tesa12345):

no. im really confused

OpenStudy (tesa12345):

@Blank

OpenStudy (kropot72):

What part of the solution do you have trouble with? The answer is given in the post by @francisgcm

OpenStudy (tesa12345):

is it a half?

OpenStudy (tesa12345):

wait no, its 2/3 right?

OpenStudy (kropot72):

Correct.